GreenHab Report – Nov 8th

Crew 214 GreenHab Report 8-11-2019

GreenHab Officer: Guy Murphy

Environmental control: Door left open for ventilation early the afternoon then reclosed early evening.

Average temperatures: –

Hours of supplemental light: Automatic system

Daily water usage for crops: Existing plantings watered by Outpost team.

Daily water usage for research and/or other purposes:

Water in Blue Tank _____ gallons

Time(s) of watering for crops: 08:30 am.

Changes to crops: N/A

Narrative: Apart from watering of the plants, 5 packets of vegetable seeds as well as plant labels were delivered today.

The seeds are identified as –

Snow Pea (Oregon Giant)

Cucumbers (Patio Snackers F1)

Pepper (Mini Bell Color Mix)

Tomato (Chocolate Cherry)

Perpetual Spinach (Green Leaf Chard)

(Botanical names were not provided). The seeds have not yet been planted.

Established plants introduced to Greenhab 03/11/2019

[3x] strawberry plants (Everbearing), in medium metal planter

[2x] small Aloe Vera, in medium metal planter, frost damaged

[1x] spicy orange thyme (thymus fragantissimus), small terra cotta pot

[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra cotta pot

[1x] Mint Mentha – Berries & Cream, small terra cotta pot

[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot

[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot

[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot

[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot

[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot

[1x] mint, grapefruit, large plastic pot

Seeds Planted in Greenhab 04/11/2019 (Large metal planting tub)

Greek Oregano – Origanum heracleoticum

Chives – Allium schoenoprasum

Sweet Marjoram – Origanum majorana

Lemon Balm – Melissa officialnis

Sage – Savia Officinalis

Thai Basil – Siam Queen

German/Winter Thyme – Thymus vulgaris

Harvest: None

Support/supplies needed: N/A

Science Report – Nov 8th

Crew 214 Science Report 8 November 2019

SOL: 12

Crew 214 – Expedition Boomerang Crew

Submitted by Science Officer Andrew Wheeler

Science Operations: No scheduled experimentation was carried out during the day. All micrometeorite samples collected during the rotation have been packaged and are awaiting dispatch to the Principle Investigator. All unused gypsum and regolith samples have been placed in storage awaiting use by Crew 215.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. Plants were watered by the Outpost crew. Additional seeds have been provided for sowing of further varieties.

Science Dome: Cleaned and tidied for Crew 215.

Closed Loop Food Waste Study: The final day of this study was concluded with the day’s waste for Sol 12 being: 28 grams; consisting of 99 calories, 2g fat, >1g saturated fat, 259mg sodium, 17g carbohydrates, 3g fibre, 3g sugar, 7g protein. For the entire 13-day rotation in sim, over 880g of food and nearly 4000 calories were wasted at a rate of approximately 7g of food per person per day. Carbohydrates contributed 63% of total calories wasted with fat and protein contributing 20% and 17% respectively.

EVA: No EVAs were performed.

EVA Suit Maintenance: Suit 10 and Suit 5 (helmet) remain unusable.

Additional Activities: N/A

Operations Report – Nov 8th

Crew 214 Operations Report 8-11-2019

SOL: 12

Name of person filing report: Andrew Wheeler

Non-nominal systems: Opportunity, Suit 10

Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning

Generator: run

Hours run: 13h 45m

From what time last night: 1745

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 76%

Diesel Reading – 55%

Station Propane Reading – 77%

Ethanol Free Gasoline: N/A

Water (loft tank): 20 gallons

Water Meter: 1456346 units

Water (static tank): 332 gallons

Static to Loft Pump used – no

Water in Green Hab: 225 gallons

Water in Science Dome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: not at hab

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Spirit rover used: no

Hours: 108.8

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 100%

Currently charging: yes

Opportunity rover used: no

Hours: 65.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: no

Curiosity rover used: no

Hours: 114.7

Beginning charge: (Before EVA) : 100%

Ending charge: (On return from EVA, before charging): 100%

Currently charging: yes

Notes on rovers: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no

Reason for use: N/A

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were used today: 0

Notes on ATVs: Assigned to the director

HabCar used and why, where? HabCar returned from service in Bicknell.

CrewCar used and why, where? CrewCar returned from service in Bicknell.

General notes and comments: N/A

Summary of internet: Functional

Summary of suits and radios: Nominal

Summary of Hab operations: All systems nominal

Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered. Additional seeds were catalogued in preparation for planting. GreenHab heater turned on automatically at 1700.

Summary of Science Dome operations: All systems nominal. Sample processing and experiments completed. All equipment used has been cleaned and returned to storage. Unused samples for ongoing experiments by Crew 215 placed in storage.

Summary of RAM operations: Not used

Summary of any observatory issues: Nominal

Summary of health and safety issues: All crew healthy

Questions, concerns and requests to Mission Support: Loft tank to be depleted to allow cleaning during crew changeover. Opportunity rover fully charged, however, the battery light is flashing and there is a bubbling sound coming from where the battery is located. It has been unplugged as a precaution.

Journalist Report – Nov 8th

Journalist Report Nov 8

Journalist: Sandy Dance

Crew 214 Sol 12 Journalist Report

by Sandy Dance

We’ve done the EVAs, we’ve done the science, its all over now bar the shouting! Well, actually the cleaning! Yes, we spent the morning vacuuming, mopping, inventoring, and general tidying up! Otherwise its all pretty quiet today.

Last night we had great success! Using the bread maker to make pizza dough, we then used the brand new pizza pan to create a remarkably genuine pizza using the dehydrated tomato, capsicum and cheese, and spam, salmon, and herbs. Probably one of our best evening meals so far!

It has been really interesting exploring how to do field science under the restrictions of wearing space suits, communicating by radio and commuting to the field by electric rover. The other important aspect is how to live in a small group under very constraining circumstances, eating a quite restricted diet of long life foods, and remaining functional. Albeit our 2 or 4 weeks is quite short compared with the 2 or 3 years of an actual mission. In summary I think what we are doing here can genuinly contribute to understanding what will be required on Mars itself!

Its been a great two weeks, we’ve learnt a lot. I’ll never look up and see Mars the same again!

Science Report Nov 8th

Crew 214 Science Report 8 November 2019

SOL: 12

Crew 214 – Expedition Boomerang Crew

Submitted by Science Officer Andrew Wheeler

Science Operations: No scheduled experimentation was carried out during the day. All micrometeorite samples collected during the rotation have been packaged and are awaiting dispatch to the Principle Investigator. All unused gypsum and regolith samples have been placed in storage awaiting use by Crew 215.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. Plants were watered by the Outpost crew. Additional seeds have been provided for sowing of further varieties.

Science Dome: Cleaned and tidied for Crew 215.

Closed Loop Food Waste Study: The final day of this study was concluded with the day’s waste for Sol 12 being: 28 grams; consisting of 99 calories, 2g fat, >1g saturated fat, 259mg sodium, 17g carbohydrates, 3g fibre, 3g sugar, 7g protein. For the entire 13-day rotation in sim, over 880g of food and nearly 4000 calories were was wasted at a rate of approximately 7g of food per person per day. Carbohydrates contributed 63% of total calories wasted with fat and protein contributing 20% and and 17% respectively.

EVA: No EVAs were performed.

EVA Suit Maintenance: Suit 10 and Suit 5 (helmet) remain unusable.

Additional Activities: N/A

Sol Summary Report Nov 8th

Sol:12

Summary Title: Final Day on Mars

Author’s name: Guy Murphy

Mission Status: Final day in full simulation mode before waking up on Earth again tomorrow morning.

Sol Activity Summary: Today was dedicated to cleaning the campus infrastructure, including both levels of the Hab, the Science Dome and the RAM. Dianne completed her food waste data collection in evening. 3 cooked meals were prepared today.

Look Ahead Plan: Changeover to Crew 215 will occur tomorrow, with 4 new crew members arriving, and 2 departing on Sunday.

Anomalies in work: None

Weather: Another sunny, clear day on Mars

Crew Physical Status: Crew in excellent health.

EVA: None undertaken

Reports to be filed: Sol Summary, Journalists Report, Greenhab Report, Science Report,

Operations Reports, Photos.

Support Requested: None